OpenStudy (anonymous):

Help?! Solve the equation: 87+2x ------- =.5 216+3x

OpenStudy (anonymous):

multiply both sides of the equation by 216+3x: 87 + 2x = 0.5 (216 + 3x) 87 + 2x = 108 + 1.5x 0.5x = 108-87 = 21 x = 21 / 0.5 = 42

OpenStudy (anonymous):

So the answer comes out to be 42?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I understand :) Thank you jimmyrep!

OpenStudy (anonymous):

as the computer said in The Hitchhikers Guide to the Galaxy

OpenStudy (anonymous):

no probs

OpenStudy (anonymous):

here we have been given that (87+2x)/(216+3x) = 5/10 multiply both sides by (216+3x) now here see we would get 87 + 2x = 5(216+3x) /10 87 + 2x = (1080+15x)/10 870 + 20x = 1080 + 15x transpose 870 to right hand side and 15x to left hand side 20x-15x = 1080-870 5x = 210 x= 42

OpenStudy (anonymous):

If the Tigers wanted to reach a winning percentage of .06, how many games would it take? (Considering they only win 2 out of 3 games.) Would you need more information to answer this question?

OpenStudy (anonymous):

*(Considering they only win 2 out of 3 games for the rest of the season.) .5 from my last question was their winning percentage.